From the information in the Data section of the textbook, calculate the equilibrium constant at 338 K for the reaction: assum step by step pls

can you only do uestion 2 and 3 for me

step by step
Answer #1

2. Reaction is:

N2O4(g) <--> 2NO2(g)
Initial (bar) 1 0
Change (bar) -\alpha +2\alpha
Equilibrium (bar) 1-\alpha 2\alpha

Kp = (PNo2)2 / PN2O4

PN2O4 = xn2O4 P = [(1-\alpha) / (1+\alpha)]P

and PNO2 = xNO2 P = [2\alpha/(1+\alpha)] P

Kp = ([2\alpha/(1+\alpha)] P)2 / ([(1-\alpha) / (1+\alpha)]P) = 4\alpha2P / (1-\alpha2)

K = 4(0.655)2(1 bar) / (1-(0.655)2) = 3.00 [not 300]

3. As we know

\DeltaG = -RT ln Kc = -(8.314 J /K.mol * 298.15 K) ln (9.18*10-8) = -40165.93 J/mol = -40.166 kJ/mol

Now,\DeltaGrxn = \DeltaG products - \DeltaGreactants

-40.166 kJ/mol = \DeltaG(AB2,(aq) ) - \DeltaG(AB2(s))

-40.166 kJ/mol = \DeltaG(AB2,(aq) ) - (-152.4kJ/mol)

\DeltaG(AB2,(aq) ) = 112.2 kJ/mol